OpenStudy (anonymous):

Determine equation of tangent line of the function f(x)=e^2x/x^2 at x=1

OpenStudy (anonymous):

\[ f'(x)=\frac{2 e^{2 x} (x-1)}{x^3}\\ f(1)=0 \] So the tangent is paralllel to the x-axis and \(f(1)=e^2\) So the the equation of the tangent is \( y=e^2\)

OpenStudy (anonymous):

Se the graph around x=1 http://www.wolframalpha.com/input/?i=plot+e%5E%282x%29%2Fx%5E2+from+x%3D.5+to+x+%3D2

OpenStudy (anonymous):

See the graph with the tangent http://www.wolframalpha.com/input/?i=plot+%28e%5E2%2Ce%5E%282x%29%2Fx%5E2+%29from+x%3D.5+to+x+%3D2

OpenStudy (anonymous):

sorry, in my post above f'(1)=0

OpenStudy (anonymous):

we have to derivate this funcion, with the chain rule and quotient rule \[f'(x)=\frac{2xe^{2x}*(x-1)}{x^4}\] \[f'(x)=\frac{2e^{2x}*(x-1)}{x^3}\] \[f'(1)=\frac{2e^{2}*(1-1)}{1^3}\] \[\boxed{\boxed{f'(1)=0}}\]